a horse in a race accelerates at the constant rate of 61 m/s2, from rest to a speed of 19.1 m/s. It then maintain that speed and covers another 8.7 m. What is the total time of the race in seconds?


1
Expert's answer
2021-10-01T12:21:42-0400

The time of the first part "t_1" can be found from acceleration definition:


"a = \dfrac{v_f-v_i}{t_1}"

where "a = 91m\/s^2" is the acceleration, "v_i = 0" is the initial speed, and "v_f = 19.1m\/s" is the final speed. Thus, obtain:


"t_1 = \dfrac{v_f}{a}"

The time of the second part "t_2" can be found from the equation of motion with constant speed:


"t_2 = \dfrac{d}{v_f}"

where "d = 8.7m" is the distance.

Finally, the total time is:


"t = t_1 + t_2\\
t = \dfrac{v_f}{a} + \dfrac{d}{v_f}\\
t = \dfrac{19.1m/s}{61m/s^2} + \dfrac{8.7m}{19.1m/s} \approx 0.77 s"

Answer. 0.77 s.